1. Cart A has a mass of 2.2kg and a velocity of 2.6. Cart B has a mass of 2.4 kg and a velocity of -2.3. What is the momentum of the two cart systems?
  2. Object A has a small mass, and Object B has a large mass. They both have the same momentum. Which one has the largest Kinetic Energy? How do you know?
  3. A 1.0 kg ball of clay moving at 2.12 m/s collides and sticks to 6.8 kg bowling ball that is initially at rest. The magnitude of the momentum of the ball and clay after the collision is:
1
Expert's answer
2020-07-08T13:15:01-0400

1)


"P=m_1v_1+m_2v_2=2.2(2.6)-2.3(2.4)=0.2\frac{kg\cdot m}{s}"

2)


"K_1=\frac{p^2}{2m_a},K_2=\frac{p^2}{2m_b}"

"\frac{K_1}{K_2}=\frac{m_b}{m_a}>1"

Object A has the largest Kinetic Energy.

3)


"P_f=P_i=m_1v_1=1(2.12)=2.12\frac{kg\cdot m}{s}"